OpenStudy (anonymous):

6y – 6 = 4y + 16 y = 2.2 y = –2.2 y = 11 y = 5 3(x – 4) = 5(x + 2) x = 11 x = –11 x = 1 x = –1

OpenStudy (allank):

The main concept here is putting like terms together. Know how to do that?

OpenStudy (anonymous):

Yes kind of-- Would it be: 6y - 4y = 2y?

OpenStudy (allank):

Not really, for the first one, you would first have 6y-4y=16+6 Know how I did that?

OpenStudy (anonymous):

OH you put the variables on one side then the constants on the other?

OpenStudy (anonymous):

Would the answer be answer choice C (for the 1st one.)

OpenStudy (allank):

Yes, that's right.

OpenStudy (anonymous):

Yay!!!!!!!!!!!!

OpenStudy (anonymous):

Now on to the next one?

OpenStudy (allank):

Sure. You'll want to 'open the brackets' first.

OpenStudy (anonymous):

Okay?

OpenStudy (allank):

That's just math slang for something like: 5(x+3) = 5x+15

OpenStudy (anonymous):

Oh. Haha-as you can see-I am not a math whiz! ;) Ok--so...?

OpenStudy (allank):

Do that for both sides of the equation. Sides, meaning the left and right of the equals sign.

OpenStudy (anonymous):

3x - 1 and 5x + 10?

OpenStudy (allank):

Almost got it. 3(x – 4) = 5(x + 2) gives you 3x-12 = 5x + 10

OpenStudy (anonymous):

Ok. So now... 12 divided by 3 and 10 divided by 5?

OpenStudy (anonymous):

okay multiply the 3 into the x and the -4

OpenStudy (allank):

Remember, put like terms together.

OpenStudy (anonymous):

then on the other side multiply the 5 into the x and the 2

OpenStudy (anonymous):

so yu get 3x-12=5x+10

OpenStudy (anonymous):

now combine the like terms. 3x and 5x, -12 and 10

OpenStudy (anonymous):

subtract the 3x from both sides and subtract 10 from both sides too.

OpenStudy (anonymous):

got it so far @JillieW

OpenStudy (anonymous):

Ok then you get 5x - -12?

OpenStudy (anonymous):

Yea I got it so far. :)

OpenStudy (anonymous):

no yu should get -22=2x

OpenStudy (anonymous):

Ok- so x =-11?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

YES! I got a 5/5. Phew. LOL---there are 3 more lessons to go. LOL

OpenStudy (anonymous):

haha have fun. hope yu learning it.

OpenStudy (anonymous):

How long are you going to be on? I have to go take my dog out for a loooong walk---and then take a shower-then I will come back and do school lol.

OpenStudy (anonymous):

How long are you going to be on Open Study?

OpenStudy (anonymous):

i'll be for like 5hours maybe.

OpenStudy (anonymous):

Ok. Well-it is like 3:01 PM here my time-so I will be back at 3:45. Thanks so much for your help again!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

Alright-I will close this Q now. Thanks again!